Circular No. 7490

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dj IN NGC 735 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), of an apparent supernova on unfiltered images taken on Sept. 8.4 (mag ∼ 17.4) and 9.4 UT (mag ∼ 17.1). The new object is located at α = 1^{h}56^{m}40^{s}.63, δ = +34^{o}10′12′′.4 (equinox 2000.0), which is 32′′.1 east and 24′′.3 south of the nucleus of NGC 735. A KAIT image of the same field taken on Aug. 25.4 showed nothing at the position of SN 2000dj (limiting mag ∼ 19.0).

CI AQUILAE S. Mazuk, R. J. Rudy, D. K. Lynch, and C. Venturini, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and T. Armstrong, Thomas K. Armstrong Development Corporation, report 0.8-2.5-μm spectrophotometry of CI Aql using the Shane 3-m telescope and the Aerospace Near Infrared Imaging Spectrograph (NIRIS) on July 18.80 UT, ∼ 74 days after peak brightness: “The excitation of the emission-line gas has increased noticeably since the May 9 infrared observations of Wilson and Dunscombe (IAUC 7426). Specifically, the infrared lines of C I and N I have disappeared, as have the fluorescently excited lines of O I. The absorption in He I 10830-nm is gone, and He II lines are now prominent. The nova has not yet entered its coronal phase, but [Si VI] at 1.9629 μm is just emerging. J, H, and K magnitudes determined from the spectrophotometry are 11.0, 10.6, and 10.1, respectively.”

V1494 AQUILAE Venturini, Rudy, Lynch, Mazuk, Puetter, and Armstrong also report

0.8–2.5-μm spectrophotometry (as above) of V1494 Aql = N Aql 1999 No. 2 on July 19.33 UT, ∼ 226 days after peak brightness: “This object is in a strong coronal phase with lines of [Si VI], [Si VII], [Ca VIII], [S VIII], and [S IX] present. The latter, at 1.2523 μm, requires photon energies > 329 eV to produce, and it is the highest excitation line that we detect. The H I, He I, and He II features have peculiar flat-topped profiles with line widths (full width at half maximum) of 2800 km/s. The J, H, and K magnitudes derived from the spectrophotometry are 9.7, 9.9, and 9.2, respectively.”

2000 September 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green